A capacitor consists of two square metal plates, each measuring 5.00x10^-2 m on a side. In between the plates is a sheet of mica measuring 1,00x10^-4 m thick. What is the capacitance of this capacitor


1
Expert's answer
2021-06-08T09:35:45-0400

The capacitance of the parallel plate capacitor


"C=\frac{\kappa \epsilon_0 A}{d}\\
=\frac{5\times 8.85\times 10^{-12}\times (5.00\times 10^{-2})^2}{1.00\times 10^{-4}}=1.11\times 10^{-9}\;\rm F"